# A Survey of Protocol Fuzzing

### v2

@@ Abstract.
Communication protocols form the bedrock of our interconnected world, yet vulnerabilities within their implementations pose significant security threats. Recent developments have seen a surge in fuzzing-based research dedicated to uncovering these vulnerabilities within protocol implementations. However, there still lacks a systematic overview of protocol fuzzing for answering the essential questions such as what the unique challenges are, how existing works solve them, etc. To bridge this gap, we conducted a comprehensive investigation of related works from both academia and industry. Our study includes a detailed summary of the specific challenges in protocol fuzzing, and provides a systematic categorization and overview of existing research efforts. Furthermore, we explore and discuss potential future research directions in protocol fuzzing. This survey serves as a foundational guideline for researchers and practitioners in the field.
@@ 1. Introduction
Communication protocols, such as TCP (Transmission Control Protocol) (Balakrishnan et al., 1995), TLS (Transport Layer Security) (Dierks, 2008), Bluetooth (SIG, 2016), etc., serve as the cornerstone of communication by defining the rules for message exchange between parties. As these protocols underpin publicly accessible services, their security is paramount and the vulnerabilities contained can lead to severe consequences. A stark illustration of this is the Heartbleed vulnerability in OpenSSL, an implementation of the TLS protocol. Upon its disclosure, Heartbleed was found to affect over 17% of servers worldwide (Synopsys, 2014; Project, 2022; Mutton, 2014), demonstrating the extensive impact a single vulnerability can have. Moreover, recent statistical analyses signal an upward trend in high-risk software vulnerabilities within network services (Hawkes, 2022), underscoring the increasing risks to network security. Given this context, the development of automated methods to detect vulnerabilities in network protocol implementations is not just beneficial but essential for the safeguarding of modern network services.
Fuzzing, as a software testing technique, was brought to the forefront by an empirical study conducted by Miller et al. in 1990 (Miller et al., 1990). This method involves the generation of a large number of random, mutated testcases aimed at triggering abnormal runtime behaviors within a software program. Due to its simplicity and scalability, fuzzing has proven to be highly effective at uncovering a wide array of bugs, leading to its widespread adoption (Zhu et al., 2022a; Manes et al., 2021; Liang et al., 2018; Li et al., 2019, 2021; Zhang et al., 2021b; Li et al., 2020; Zheng et al., 2022; Zhang et al., 2023c; Du and Li, 2023; Zhang et al., 2021a, 2023a). However, fuzzing protocol implementations, as opposed to general software like command-line tools (Mathis et al., 2019; Pham et al., 2019; Wang et al., 2017), introduces additional challenges. These complexities are largely due to the peculiarities associated with effectively testing the intricate communication logic that protocols entail, ranging from methodological considerations to tool-specific requirements. In response to these challenges, there is a notable trend towards creating advanced fuzzing methods tailored explicitly for protocol testing (Beurdouche et al., 2015; de Ruiter and Poll, 2015; Somorovsky, 2016; Pham et al., 2020; Meng et al., 2024; Luo et al., 2023; Garbelini et al., 2022a, c; Ba et al., 2022). Despite this progress, there still remains a significant gap in research dedicated to systematically examining the distinctive challenges inherent to this field, thoroughly summarizing the existing solutions and discussing future directions. To fill this gap, we extensively discussed and analyzed protocol fuzzing specifics in the following content of this article.
@@ 1.1. Motivation
Protocols are the essential rules that dictate how our devices and applications communicate, making them both pervasive and critically important. Because these protocols are everywhere, ensuring they are secure against potential threats is of utmost importance. Fuzzing plays a key role in finding and fixing security issues within these systems. In light of this, building the first end-to-end guide covering the overview and specifics of protocol fuzzing is highly valuable for both researchers and those in the tech industry.
Protocol fuzzing presents unique challenges that set it apart from general application fuzzing, grounded in the intricacies of the communication protocols themselves. Firstly, there’s the need to adhere to strict rules that dictate not just the structure of the messages but also the strict sequence and context in which these messages are sent and received (802, 2021, 2020; Dierks, 2008; Rescorla, 2012). This makes the testing process complex as it requires an in-depth grasp of how these communication protocols operate and change over time. Secondly, protocols are built to address various attributes beyond simple message exchange. They must account for factors like timing and how multiple messages or actions can happen simultaneously, which introduces more variables into the mix when testing for security issues (Jelassi et al., 2012; Katz and Shin, 2006; Jung et al., 2001). Thirdly, the widespread use of protocols across different technology levels and systems adds another layer of complexity. They are embedded everywhere, from hardware up to the software applications we interact with daily, leading to diverse testing scenarios and discovering potential vulnerabilities in every layer (Ruge et al., 2020; Wu et al., 2021; Garbelini et al., 2020, 2022a, 2022c; Fang et al., 2021; Luo et al., 2020). Given these realities, it becomes imperative to establish a comprehensive understanding of protocol-specific challenges.
Many protocol fuzzing works have been completed but no systematic review on protocol fuzzing has been conducted thus far. Although some survey articles (Liang et al., 2018; Manes et al., 2021; Zhu et al., 2022b) about traditional software fuzzing are available, they cannot provide a systematic overview of the current status and future directions based on existing works solving protocol-specific challenges.
1.2. Research Questions
This survey aims to provide an overview of the protocol-specific challenges, the corresponding solutions, and the future directions. Specifically, this survey answers the following questions:
RQ1: What are the differences between traditional fuzzing and protocol fuzzing?
RQ2: How do existing works address the challenges in protocol fuzzing?
RQ3: What are the potential future directions?
@@ 1.3. Collection Strategy
In this survey, we focus on stateful network protocols and the various techniques that are directly related to the fuzz testing of their implementations. To collect the relevant publications, we followed the procedures depicted in Fig. 2. First, we performed a preliminary reading and summarized 12 different keyword combinations that can be used to search related works. Then, searching these keyword combinations in Google Scholar, we collected 535 publications published from 2013 to 2023. After that, we manually filtered out the papers irrelevant to protocol fuzzing or not published in the influential publications listed in Table 1. At this time, the number of papers was reduced to 78. Note that all preprint papers were kept to remove publication bias (Yasin et al., 2020). And a paper is relevant if its key contribution is in the scope of protocol fuzzing or that paper is a bug detection tool and has picked at least one protocol implementation as its evaluation target. The latter criteria is based on the heuristic that likely a bug detection tool have proposed protocol specific techniques if it uses protocol implementations as its evaluation targets. Next, we performed snowballing and inverse snowballing to obtain a more comprehensive view. Six more papers were added in this procedure. Finally, we applied the above collection process to the released talks of several mainstream industrial security conferences such as BlackHat. 21 industrial works were added, including 18 related talks and three open-source protocol fuzzers with more than 50 stars on Github. The ascending trajectory of publications, as illustrated in Fig. 2, underscores the burgeoning research interest in protocol fuzz testing, affirming its emergence as a focal point within the field.
The rest of the paper is organized as follows. Section 2 introduces the background knowledge of protocol fuzzing. Section 3 introduces the main differences between general fuzzing targets and protocols, then summarize the major enhancements of existing protocol fuzzers. The next three sections detail the existing techniques for each key component of protocol fuzzing. Section 4 discusses the progress in input generator component. Section 5 introduces the techniques for improving the executor component. Section 6 manifests the taxonomy of oracles used in the bug detector component. Section 7 offers future directions.
@@ 2.1. Communication Protocols
A communication protocol is a set of rules that enables the exchange of information between two or more entities within a communication system, utilizing any form of physical quantity variation. The implementation of a communication protocol generally involves multiple phases (Comer, 2013). First, the protocol is conceptually designed, which includes defining the rules, behaviors, and functions it will perform based on the protocol’s needs, taking into account factors such as efficiency, reliability, scalability, and security. The outcome of the design phase is a specifications. Then, during the development phase, the protocol design is translated into concrete implementations. This can be in the form of software, hardware or a combination of both. Once developed, the protocol undergoes rigorous testing to confirm that it adheres to the protocol specifications and meets performance and reliability requirements. Among them, fuzzing, which this paper focuses on, is a commonly used technology for testing protocol implementations. Eventually, the protocol implementation will be deployed in a real-world environment.
In addition to the fundamental task of data exchange, protocols encompass a myriad of other functionalities critical to communication, introducing new layers of complexity (Marsden, 1986). This includes tasks such as routing, detection of transmission errors, managing timeouts and retries, confirmations, flow control, and sequence control. As a typical example, TCP (Transmission Control Protocol) (Balakrishnan et al., 1995) contains the following functionalities for guaranteeing optimal communication:
Acknowledgment: Acknowledgment is a mechanism to confirm the receipt of data packets. This process is crucial for ensuring the reliability of data transmission, as it allows the sender to know whether the data has reached its intended destination. In TCP, when a data packet is received, the recipient sends back an acknowledgment to the sender to confirm receipt.
Sequence control: Sequence control ensures that data packets are received and processed in the order they were sent. TCP segments are sequenced with a sequence number.
Error handling: Error handling involves detecting and correcting errors that occur during data transmission. TCP includes error-checking features. Every TCP segment contains a checksum field, which is used to check for data integrity. If a segment is found to be corrupted (i.e., the data does not match the checksum), it is discarded, and TCP will handle retransmission.
Each of these functionalities embodies a set of strategies and implementations that collectively ensure the efficacy and reliability of communication protocols. The intricate integration of these functions demonstrates the sophisticated nature of protocol design and its pivotal role in modern communication systems.
@@ 2.2. Types of Protocols
Protocols can be classified from various perspectives, such as functionality, the accessibility of their specifications, and their alignment with the layers of the OSI network reference model.
From a functional standpoint, protocols exhibit a broad spectrum of varieties, each tailored to fulfill unique operational objectives. For instance, security protocols are primarily designed to ensure the integrity and confidentiality of data, exemplified by TLS (Dierks, 2008) and DTLS (Datagram Transport Layer Security) (Rescorla, 2012). Routing protocols, such as BGP (Border Gateway Protocol), are dedicated to efficiently managing the routes that data packets traverse across the network. Furthermore, application protocols, like HTTP (Hypertext Transfer Protocol) for web services and SMTP (Simple Mail Transfer Protocol) for email, are specialized to enable specific functionalities at the application layer.
When considering the availability of protocol specifications, a distinction is drawn between open protocols and proprietary protocols. Open protocols, like TCP, have publicly accessible specifications, allowing for widespread scrutiny and implementation. In contrast, proprietary protocols such as Microsoft’s RDP (Remote Desktop Protocol) (Microsoft, 2007), are governed by individual entities, with specifications that are not fully public. The availability of protocol specifications is crucial for various stages of fuzzing, such as crafting inputs, constructing state machines, and detecting bugs. It is important to clarify that the classification into open and proprietary pertains solely to the availability of specifications, and is independent of the accessibility of source codes for protocol implementations.
Regarding the statefulness, protocols are bifurcated into stateful and stateless categories. Stateful protocols, such as TLS (Somorovsky, 2016; de Ruiter and Poll, 2015) and TCP (Jero et al., 2018), necessitate multiple interaction rounds. Stateless protocols, like UDP and HTTP, do not maintain state information across requests.
Based on the OSI network reference model, protocols can be classified into seven distinct layers: physical, data link, network, transport, session, presentation, and application. The protocol layers each solve a distinct class of communication problems. Among them, the lower-level protocols have higher coupling with physical hardware. It is pertinent to note that not every protocol aligns precisely with a single layer in the OSI model. For example, TLS/DTLS contains the functionality of the session and representation layers; the Wi-Fi protocol contains the main functionality of the physical and data link layers (802, 2021; Chen et al., 2005). Given varying interpretations of protocol layering in numerous sources, we categorize these protocols based on their primary functions.
@@ 3.1. Differences between protocol fuzzing and traditional fuzzing
In this subsection, we delve into the unique challenges associated with protocol fuzzing as identified in the literature, addressing RQ1. Table 2 encapsulates two primary distinctions between protocol implementations and general fuzzing targets. These differences not only highlight the specificities of protocol fuzzing but also correspond to a set of inherent challenges.
| General Fuzzing Targets | Protocol Implementations | |
| Communication Complexity | Low | High |
| Testing Environment | Unconstrained | Constrained |
@@ 3.1.1. High communication complexity
Respecting Semantic Constraints In Communication. Protocols serve as the backbone for facilitating communication between different systems by providing a standardized set of rules for message exchange. This communication is inherently complex, often involving a multi-round process where multiple steps must be sequentially executed for the exchange to be successful. Such protocols inherently demand implementations that are stateful, with each stage of communication building upon the previous one (802, 2021, 2020; Dierks, 2008; Rescorla, 2012). In testing scenarios, this means that deeper layers of the protocol implementation cannot be tested until the earlier constraints are satisfactorily met – these are the strict semantic constraints inherent in communication protocols. Semantic constraints come in two primary forms: intra-message and inter-message constraints. Intra-message constraints pertain to the structure and content of individual messages, ensuring that data fields are syntactically correct and semantically meaningful within the context of that message. Taking TCP as an example, in a TCP segment, there are several critical fields such as the source port, destination port, sequence number, acknowledgment number, data offset, and control flags (like SYN, ACK) (Marsden, 1986). Each of these fields must adhere to specific formats and rules. Inter-message constraints, on the other hand, govern the relationship and sequence of multiple messages, requiring that they adhere to the established protocol sequence and context for the conversation to progress (Comer, 2013). For instance, the establishment of a TCP connection involves a “three-way handshake” process: the client first sends a SYN message, followed by the server responding with a SYN-ACK message, and finally, the client sends an ACK message to complete the connection. Violations of either type of constraints during communication can result in the fuzzing nonprogressive (Pham et al., 2020; Somorovsky, 2016; Zou et al., 2021; Jero et al., 2018; Vanhoef, 2017).
@@
Testing Different Properties of Communication Process. Besides basic message exchange functionality, protocols need to guarantee a series of extra features that forming a more secure or reliable communication such as timing requirements, authentication, confidentiality, and concurrency(Jelassi et al., 2012; Katz and Shin, 2006; Jung et al., 2001). Effectively testing these attributes in the implementation requires a more complex form of testing that goes beyond typical application fuzzing which mainly focuses on altering structured inputs to find issues (Zhu et al., 2022b; Manes et al., 2021). Each attribute may require the significant modification or even redesign of the fuzzing framework, including the developments of specialized input generator, feedback mechanisms, and oracles to facilitate effective testing (Hussain et al., 2021; Ji and Xu, 2023; Somorovsky, 2016; Zou et al., 2021; Vanhoef, 2017; Shi et al., 2019; Garbelini et al., 2020; Meng et al., 2021; Chen et al., 2022; Ma et al., 2023). For instance, in the context of crafting a fuzzer aiming at detecting traffic amplification attacks within protocol implementations (Krupp et al., 2022), an oracle is needed to identify disproportional request-to-response data volume ratios, indicative of an amplification factor. Concurrently, input generator needs to be adeptly redesigned to generate specific variations of protocol messages that can maximize the potential amplification factor. Moreover, the amplification factor can be used as a feedback to further enhance the search performance of fuzzing.

### v3

@@ Abstract.
Communication protocols form the bedrock of our interconnected world, yet vulnerabilities within their implementations pose significant security threats. Recent developments have seen a surge in fuzzing-based research dedicated to uncovering these vulnerabilities within protocol implementations. However, there still lacks a systematic overview of protocol fuzzing for answering the essential questions such as what the unique challenges are, how existing works solve them, etc. To bridge this gap, we conducted a comprehensive investigation of related works from both academia and industry. Our study includes a detailed summary of the specific challenges in protocol fuzzing and provides a systematic categorization and overview of existing research efforts. Furthermore, we explore and discuss potential future research directions in protocol fuzzing.
@@ 1. Introduction
Communication protocols, such as TCP (Transmission Control Protocol) (Balakrishnan et al., 1995), TLS (Transport Layer Security) (Dierks, 2008), Bluetooth (SIG, 2016), etc., serve as the cornerstone of communication by defining the rules for message exchange between parties. As these protocols underpin publicly accessible services, their security is paramount, and the vulnerabilities contained can lead to severe consequences. A stark illustration of this is the Heartbleed vulnerability in OpenSSL, an implementation of the TLS protocol. Upon its disclosure, Heartbleed was found to affect over 17% of servers worldwide (Synopsys, 2014; Project, 2022; Mutton, 2014), demonstrating the extensive impact that a single vulnerability can have. Moreover, recent statistical analyzes signal an upward trend in high-risk software vulnerabilities within network services (Hawkes, 2022), underscoring the increasing risks to network security. Given this context, the development of automated methods to detect vulnerabilities in network protocol implementations is not just beneficial but essential for the safeguarding of modern network services.
Fuzzing, as a software testing technique, was brought to the forefront by an empirical study conducted by Miller et al. in 1990 (Miller et al., 1990). This method involves the generation of a large number of random mutated testcases aimed at triggering abnormal runtime behaviors within a software program. Due to its simplicity and scalability, fuzzing has proven to be highly effective in uncovering a wide variety of bugs, leading to its widespread adoption (Zhu et al., 2022; Manes et al., 2021; Liang et al., 2018; Zhang et al., 2023b; Zheng et al., 2022; Zhang et al., 2021). However, fuzzing protocol implementations, as opposed to general software such as command-line tools (Mathis et al., 2019; Pham et al., 2019; Wang et al., 2017), introduces additional challenges. These complexities are largely due to the peculiarities associated with effectively testing the intricate communication logic that protocols entail, ranging from methodological considerations to tool-specific requirements. In response to these challenges, there is a notable trend toward the creation of advanced fuzzing methods explicitly tailored for protocol testing (Beurdouche et al., 2015; de Ruiter and Poll, 2015; Somorovsky, 2016; Pham et al., 2020; Meng et al., 2024; Luo et al., 2023; Garbelini et al., 2022a, c; Ba et al., 2022). Despite this progress, there still remains a significant gap in research dedicated to systematically examining the distinctive challenges inherent to this field, thoroughly summarizing the existing solutions and discussing future directions. To fill this gap, we extensively discuss and analyze protocol fuzzing specifics in the following content of this article.
@@ 1.1. Motivation
Protocols are the essential rules that dictate how our devices and applications communicate, making them both ubiquitous and critically important. Because these protocols are everywhere, it is of the utmost importance to ensure they are secure against potential threats. Fuzzing plays a key role in finding and fixing security issues within these systems. In light of this, building the first end-to-end guide covering the overview and specifics of protocol fuzzing is highly valuable for both researchers and those in the tech industry.
Protocol fuzzing presents unique challenges that set it apart from general application fuzzing, grounded in the intricacies of the communication protocols themselves. Firstly, there is the need to adhere to strict rules that dictate not just the structure of the messages but also the strict sequence and context in which these messages are sent and received (802, 2020; Dierks, 2008; Rescorla, 2012). This makes the testing process complex as it requires an in-depth understanding of how these communication protocols operate and change over time. Secondly, protocols are built to address various attributes beyond simple message exchange. They must account for factors such as timing and how multiple messages or actions can occur simultaneously, which introduces more variables into the mix when testing for security issues (Jelassi et al., 2012; Katz and Shin, 2006; Jung et al., 2001). Third, the widespread use of protocols across different technology levels and systems adds another layer of complexity. They are embedded everywhere, from hardware to software applications that we interact with daily, leading to various testing scenarios and discovering potential vulnerabilities at every layer (Ruge et al., [n. d.]; Wu et al., 2021; Garbelini et al., 2020, 2022a, 2022c; Fang et al., 2021; Luo et al., 2020). Given these realities, it becomes imperative to establish a comprehensive understanding of protocol-specific challenges.
Many protocol fuzzing works have been completed, but so far no systematic review on protocol fuzzing has been conducted. Although some survey articles (Liang et al., 2018; Manes et al., 2021; Zhu et al., 2022) on traditional software fuzzing are available, they cannot provide a systematic overview of the current state and future directions based on existing work solving protocol-specific challenges.
1.2. Survey Dimensions
This survey aims to provide an overview of the protocol-specific challenges, the corresponding solutions, and the future directions. Specifically, this survey is organized around the following dimensions:
Dimension 1: Examine the differences between traditional fuzzing and protocol fuzzing.
Dimension 2: Review how existing works address the challenges in protocol fuzzing.
Dimension 3: Explore potential future directions in the field.
In Section 3, we provide an in-depth examination of the distinctive differences between protocols and traditional fuzzing targets to address Dimension 1. Then, in Sections 4 to 6, we detail the techniques used in existing protocol fuzzers to address Dimension 2. Lastly, Dimension 3 is discussed in Section 7.
@@ 1.3. Collection Strategy
In this survey, we focus on stateful network protocols and the various techniques that are directly related to the fuzz testing of their implementations. To collect the relevant publications, we followed the procedures depicted in Fig. 2. First, we performed a preliminary reading and summarized 12 different keyword combinations that can be used to search for related works. Then, searching for these keyword combinations on Google Scholar, we collected 544 published articles from January 2013 to June 2024. After that, we manually filtered out papers irrelevant to protocol fuzzing or not published in influential publications listed in Table 1. At that time, the number of articles was reduced to 87. Note that all pre-print papers were kept to remove publication bias (Yasin et al., 2020). And a paper is relevant if its key contribution is in the scope of protocol fuzzing or that paper is a bug detection tool and has picked at least one protocol implementation as its evaluation target. The latter criterion is based on the heuristic that a bug detection tool probably has proposed protocol-specific techniques if it uses protocol implementations as its evaluation targets. Next, we performed snowballing and inverse snowballing to obtain a more comprehensive view. Six more papers were added in this procedure. Finally, we applied the above collection process to the released talks of several mainstream industrial security conferences such as BlackHat. 22 industrial works were added, including 20 related talks and three open source protocol fuzzers with more than 50 stars on Github. The ascending trajectory of publications, as illustrated in Fig. 2, underscores the burgeoning research interest in protocol fuzz testing, affirming its emergence as a focal point within the field.
The remainder of the paper is organized as follows. Section 2 introduces the background knowledge of protocol fuzzing. Section 3 introduces the main differences between general fuzzing targets and protocols and then summarizes the major enhancements of existing protocol fuzzers. The next three sections detail the existing techniques for each key component of protocol fuzzing. Section 4 discusses the progress in input generator component. Section 5 introduces the techniques for improving the executor component. Section 6 manifests the taxonomy of oracles used in the bug detector component. Section 7 offers future directions.
@@ 2.1. Communication Protocols
A communication protocol is a set of rules that enables the exchange of information between two or more entities within a communication system, utilizing any form of physical quantity variation. The implementation of a communication protocol generally involves multiple phases (Comer, 2013). First, the protocol is conceptually designed, which includes defining the rules, behaviors, and functions it will perform based on the protocol’s needs, taking into account factors such as efficiency, reliability, scalability, and security. The outcome of the design phase is a set of specifications. Then, during the development phase, the design of the protocol is translated into concrete implementations. This can be in the form of software, hardware, or a combination of both. Once developed, the protocol undergoes rigorous testing to confirm that it adheres to the protocol specifications and meets the performance and reliability requirements. Among them, fuzzing, which this paper focuses on, is a commonly used technology for testing protocol implementations. Eventually, the protocol implementation will be deployed in a real-world environment.
In addition to the fundamental task of data exchange, protocols encompass a myriad of other critical communication functionalities, introducing new layers of complexity (Marsden, 1986). This includes tasks such as routing, detection of transmission errors, managing timeouts and retries, confirmations, flow control, and sequence control. Each of these functionalities embodies a set of strategies and implementations that collectively ensure the efficacy and reliability of communication protocols. The intricate integration of these functions demonstrates the sophisticated nature of protocol design and its pivotal role in modern communication systems.
@@ 2.2. Types of Protocols
Protocols can be classified from various perspectives, such as functionality, the accessibility of their specifications, and their alignment with the layers of the OSI network reference model, as shown in Fig. 3.
From a functional point of view, protocols exhibit a broad spectrum of varieties, each tailored to fulfill unique operational objectives. For example, security protocols are designed primarily to ensure the integrity and confidentiality of data, as exemplified by TLS (Dierks, 2008) and DTLS (Datagram Transport Layer Security) (Rescorla, 2012). Routing protocols, such as BGP (Border Gateway Protocol), are dedicated to efficiently managing the routes of data packets traversing the network. Furthermore, application protocols, such as HTTP (Hypertext Transfer Protocol) for web services and SMTP (Simple Mail Transfer Protocol) for email, are specialized to enable specific functionalities at the application layer.
When considering the availability of protocol specifications, a distinction is drawn between open protocols and proprietary protocols. Open protocols, like TCP, have publicly accessible specifications, allowing widespread scrutiny and implementation. In contrast, proprietary protocols such as Microsoft’s RDP (Remote Desktop Protocol) (Microsoft, 2007), are governed by individual entities, with specifications that are not fully public. The availability of protocol specifications is crucial for various stages of fuzzing, such as crafting inputs, constructing state machines, and detecting bugs. It is important to clarify that the classification into open and proprietary pertains solely to the availability of specifications and is independent of the accessibility of source codes for protocol implementations.
Regarding the statefulness, protocols are bifurcated into stateful and stateless categories. Stateful protocols, such as TLS (Somorovsky, 2016; de Ruiter and Poll, 2015) and TCP (Jero et al., 2018), require multiple rounds of interaction. Stateless protocols, such as UDP and HTTP, do not maintain state information across requests.
Based on the OSI network reference model, protocols can be classified into seven distinct layers: physical, data link, network, transport, session, presentation, and application. The protocol layers each solve a distinct class of communication problems. Among them, the lower-level protocols have a higher coupling with the physical hardware. It is pertinent to note that not every protocol aligns precisely with a single layer in the OSI model. For example, TLS/DTLS contains the functionality of the session and representation layers; the Wi-Fi protocol contains the main functionality of the physical and data link layers (802, 2021; Chen et al., 2005). Given varying interpretations of the protocol layering in numerous sources, we categorize these protocols based on their primary functions.
@@ 3.1. Differences between protocol fuzzing and traditional fuzzing
In this subsection, we discuss the unique challenges associated with protocol fuzzing as identified in the literature, addressing Survey Dimension 1. Protocol implementations differ from traditional fuzzing targets in two key aspects: first, they exhibit higher communication complexity; second, their testing environments are relatively more constrained. These differences not only highlight the specificity of protocol fuzzing, but also correspond to a set of inherent challenges.
@@ 3.1.1. High communication complexity
Respecting Semantic Constraints in Communication. Protocols serve as the backbone for facilitating communication between different systems by providing a standardized set of rules for message exchange. This communication is inherently complex, often involving a multi-round process where multiple steps must be sequentially executed for the exchange to be successful. Such protocols inherently require stateful implementations, with each stage of communication building upon the previous one (802, 2021, 2020; Dierks, 2008; Rescorla, 2012). In testing scenarios, this means that deeper layers of the protocol implementation cannot be tested until the earlier constraints are satisfactorily met – these are the strict semantic constraints inherent in communication protocols. Semantic constraints come in two primary forms: intra-message and inter-message constraints. Intra-message constraints pertain to the structure and content of individual messages, ensuring that data fields are syntactically correct and semantically meaningful within the context of that message. Taking TCP as an example, in a TCP segment, there are several critical fields such as the source port, destination port, sequence number, acknowledgment number, data offset, and control flags (like SYN, ACK) (Marsden, 1986). Each of these fields must adhere to specific formats and rules. Inter-message constraints, on the other hand, govern the relationship and sequence of multiple messages, requiring that they adhere to the established protocol sequence and context for the conversation to progress (Comer, 2013). For instance, the establishment of a TCP connection involves a “three-way handshake” process: the client first sends a SYN message, followed by the server responding with a SYN-ACK message, and finally, the client sends an ACK message to complete the connection. Violations of either type of constraints during communication can result in the fuzzing becoming non-progressive (Pham et al., 2020; Somorovsky, 2016; Zou et al., [n. d.]; Jero et al., 2018; Vanhoef, 2017).
@@
Testing Different Properties of Communication Process. Besides basic message exchange functionality, protocols need to guarantee a series of additional features that form a more secure or reliable communication such as time requirements, authentication, confidentiality, and concurrency(Jelassi et al., 2012; Katz and Shin, 2006; Jung et al., 2001). Effectively testing these attributes in implementations requires a more complex form of testing that goes beyond typical application fuzzing, which mainly focuses on altering structured inputs to find issues (Zhu et al., 2022; Manes et al., 2021). Each attribute may require significant modification or even a redesign of the fuzzing framework, including the development of specialized input generator, feedback mechanisms, and oracles to facilitate effective testing (Hussain et al., 2021; Ji and Xu, 2023; Somorovsky, 2016; Zou et al., [n. d.]; Vanhoef, 2017; Shi et al., 2019; Garbelini et al., 2020; Meng et al., 2021; Chen et al., 2023; Ma et al., 2023a). For example, in the context of crafting a fuzzer that aims to detect traffic amplification attacks within protocol implementations (Krupp et al., 2022), an oracle is needed to identify disproportional request-to-response data volume ratios, indicative of an amplification factor. Currently, the input generator needs to be adeptly redesigned to generate specific variations of protocol messages that can maximize the potential amplification factor. Moreover, the amplification factor can be used as a feedback to guide the fuzzer in exploring the input space more effectively.